Your answer is partially correct. Try again. At the beginning of the year, Morales Company had total assets of $809,000 and total liabilities of $556,000. (Treat each item independently.) (a) If total assets increased $151,000 during the year and total liabilities decreased $82,000, what is the amount of stockholders equity at the end of the year? Stockholders equity 486000 (b) During the year, total liabilities increased$90,000 and stockholders equity decreased $73,000. What is the amount of total assets at the end of the year? Total assets 179747 If total assets decreased $89,000 and stockholders equity increased $113,000 during the year, what is the amount of total liabilities at the end Total liabilities 607000please help i tried every answer
Brief Exercise I- (b) During the year, total liabilities increased $90,000 and stockholders equity decreased $73,000. what is the amount of total assets at e d Exercise- ー Exercise 1 Total assets 19000 If total assets decreased $89,000 and stockholders equity incressed $113,000 during the yeer, whet is the amount of total labilities at the end of the year Review Review Results by Study Objectivs Total liabilities Click if you would like to Show Work for this questions Qpen Show Work
Answer #1
Stockholders' Equity at beginning 253000 =809000-556000
a
Stockholders' Equity 486000 =253000+151000+82000
b
Total Assets 826000 =809000+90000-73000
c
Total Liabilities 354000 =556000-89000-113000